OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (anonymous):

-5 5 9 -9

OpenStudy (imstuck):

You are solving for x, I take it?

OpenStudy (anonymous):

@iGreen

OpenStudy (imstuck):

First step is to distribute the -4 into the parenthesis and then on the other side, distribute the -6 into the parenthesis.

OpenStudy (igreen):

Distribute both -4 and -6 into the parentheses.

OpenStudy (igreen):

-4 * 2x = ? -4 * -1 = ? -6 * x = ? -6 * 2 = ?

OpenStudy (igreen):

Do you know the answers to those?

OpenStudy (igreen):

@_angeliquexoxo

OpenStudy (igreen):

-4 * 2x = -8x -4 * -1 = 4 -6 * x = -6x -6 * 2 = -12

OpenStudy (igreen):

So we have: -8x + 4 = -6x - 12 - 2 Simplify: -8x + 4 = -6x - 14 Subtract the four: -8x = -6x - 18 Add the 6x: -2x = -18 Divide: x = 9

OpenStudy (igreen):

@_angeliquexoxo